16: A fluorinated organic gas in a cylinder is compressed from an initial volume of 950 mL at 158 Pa to 494 mL at the same temperature. What is the final pressure? Answer in units of Pa.
17: The pressure on a gas at −51◦C is doubled, but its volume is held constant. What will the final temperature be in degrees Celsius? Answer in units of ◦C.
18: A sample of a gas occupies 1810 mL at 319◦C and 740 torr. At what temperature would it occupy 1586.8 mL if the pressure is kept constant? Answer in units of ◦C.
19: A meteorological balloon contains 124 L of He at 26◦C and 0.955 atm. It will burst if the volume reaches 358 L. At what temperature will the balloon burst if its pressure at that bursting point is 0.497 atm? Answer in units of ◦C.
1
Expert's answer
2020-04-19T07:09:09-0400


16). V1=950ml

P1=158pa

V2=494ml

P2=?

"P1V1=P2V2"

P2= (158*950)/494

P2=303.85Pa

17). T1=222.15K

P1= 1

P2= 2

T2=?

"P1T2=P2T1"

T2=(2*222.15)1

T2=444.3K

T2= 171.15OC

18). V1=1810ml

T1=592.15K

P= 740= CONSTANT

V2=1586.8ml

T2=?

"V1T2=V2T1"

T2=(1586.8*592.15)/1810=519.13K

T2=245.980C

19). 124 L of He at 26◦C and 0.955 atm

V1=124L

T1=299.15K

P1=0.955atm

P2=0.497

V2=358L

T2=?

"P1V1T2=P2V2T1"

T2= (0.497*358*299.15)/(0.955*124)

T2=300.73K

T2=27.580C